Crew 263 Journalist Report Mar1st

category journalist-report]

Sol 8 – Time for geology!

Author: Nicolas Wattelle

I think we are now well settled in the MDRS routine. Our mission plan is driving our daily life, and it feels comfortable to rely on it as we are now accustomed to the latter. See, even the (firstly “annoying”) morning sport sessions are becoming a formality to us! Or almost, I may exaggerate…

We woke up like every morning at 6:45am, to be ready for the 7 am session. This week, we begin a study of our circadian cycle. We have here two “ArmBands” given by CNES. These devices allow us to determine, by wearing them during 48 hours, our “body” cycle. When we talk about circadian, we talk about all that is cyclical in our physiology. The energy expenditure, the intensity of activity, the body temperature… How is this cycle impacted by our life on Mars? We will see! Today, Valentine and Mathéo started to wear them just before working out, and so for the two coming days. We will all wear one at some point during this mission.

Then, we continued our morning routine, as everyday: breakfast (maybe there I can add that we have more and more deep conversations), preparation for the EVAs, then separation of the crew for their respective experiments. But today’s EVA was special and has been excitingly awaited: the first for the geological study thanks to the handheld LIBS (Laser Induced Breakdown Spectroscopy) analyzer from SciAps, while using the Hololens to add an Artificial Intelligence support to help the working astronauts! For this one, Cerise, Marine, Elena and Léa went out, and headed to the “Stream Bed Connector” zone. To be honest, this one asked more logistic than the previous ones, it thus began a quarter late to be sure everything was fine before going out. And for a first try it went quite well and it is very promising for the later EVAs. During the EVA, the “laser gun” stopped functioning, but it was not a problem since the martionauts brought back 3 precisely chosen samples, and we were able to turn the device on again in the station. While I am writing these words, the two freshly new geologists of the crew are processing the data taken out of these samples with lots of interest! It is very nice to see this study begin!

I asked a new question to my teammates:

According to you, and at this point of our mission, how do you feel about living isolated?

This question divided the crew on the first thought. The first « team », made of Elena, Marine and I, first mentioned the comfort of leaving behind our daily mental load. I think this aspect is notable as to live a faithful simulation (and we do!) we have to leave some of our terrestrial workload behind. The others more quickly pointed out the fact that they sometimes miss their close ones, especially when we left them in a particular context in which we could be needing news from them. This is kind of weird as said Valentine, we live cut out from the world, but we live really close to each other: we live alone, but together.

PS: I cannot end this report without thanking our crew engineer Mathéo, who always takes advantage of his not too busy days (like today) to cook some bread and wonderfully delights for the whole crew. This thing is beginning to look like a tradition. We are all fans of the delicious smell in the Hab when we all gather to write our reports!

Crew 263 EVA Report Mar 1st

category eva-report]

Crew 263 EVA Report 01-03-2022

EVA # 12

Author: Marine Prunier

Purpose of EVA: Analysis of geology transition zones using a handheld LIBS (Laser-Induced Breakdown Spectroscopy) analyzer.

Start time: 10:14

End time: 11:46

Narrative: EVA Participants entered the airlock at 10:14. Depressurization ended at 10:19. EVA participants checked the rover batteries. The obtained information is:

Perseverance: Usage Duration – 228.1 Battery Charge – 100%

Spirit: Usage Duration – 157.2 Battery Charge – 100%

Opportunity: Usage Duration – 84.7 Battery Charge – 100%

Curiosity: Usage Duration – 168.7 Battery Charge – 100%

The EVA participants took Opportunity and Curiosity at 10:25 and drove on Entrance Road until Cow Dung Road. They drove on Cow Dung Road until Marble Ritual, after Marble Ritual the road was partially flooded on 1-2 meters, they stopped and parked the rovers along Cow Dung Road at 10:29. The EVA participant walked on Cow Dung Road until the junction with Stream Bed Connector (10:38).
They walked on Stream Bed Connector and reached two points of interest at 10:45 (the two points of interest were located 20 meters from each other, GPS location 519297 E 4251286 N). The EVA participants stayed about 35 minutes in this area, they proceeded on sketches and pictures. A software issue with the LIBS analyzer prevented the participants to use it on the field, therefore three samples of rocks have been taken for further analysis at the station.

At 11:19, the participants left the point of interest. At 11:39 they reached the rovers parked on Cow Dung Road. When starting the return to the station, a tourist car surprised the participants from behind, they waited for them to evacuate the road and returned safely to the Hab.

At 11:41, the participants arrived at the Hab:

Opportunity: Usage Duration – 84.8 Battery Charge – 99%

Curiosity: Usage Duration – 168.9 Battery Charge – 97%

They entered Airlock and at 11:46 the pressurization was over.

Destination: Zone F -final point of interest reached located on the Stream Bed Connector road at 12S 519297 E 4251286 N.

Participants: Marine Prunier (EVA leader), Léa Rouverand, Cerise Cuny, Elena Lopez-Contreras. HabCom was Valentine Bourgeois.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road, Stream Bed Connector Road.

Mode of travel: Rover on Entrance Road and Cow Dung Road till Marble ritual, foot on Cow Dung Road and on Stream Bed Connector.

Operations Report – February 27th

Crew 263 Operations Report 27-02-2022

SOL: 6

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Opportunity rover used: No

Hours: 84.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Curiosity rover used: No

Hours: 168.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Shannon

General notes and comments: NA

Summary of Hab operations:

WATER USE: 13 gallons

Water (static tank): 417 gallons

Water (loft tank): 42 gallons

Water Meter: 01561467 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater: On

Toilet tank emptied: No

Summary of internet: NA

Summary of suits and radios: The microphone of an earpiece is not working anymore, it would be great to have a new one if possible. Thank you!

Summary of GreenHab operations: NA

Summary of ScienceDome operations:

Dual split: Off

SOC: Shannon worked all day to fix the SOC, the generator is working correctly except that the SOC is not calculating correctly.

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO Report

Questions, concerns and requests to Mission Support: NA

Supplemental Operation – February 25th

Supplemental Operations Report 25Feb2022

Name of person filing report: Shannon Rupert

Reason for Report: Changing of the guard

Non-nominal systems: GreenHab

Action taken for non-nominal systems: All crops were removed from the GreenHab and it will be shut down quite possibly for the rest of the field season

Generator: Still being manually run at night. 967.5 hours

ScienceDome Dual Split: Off

Solar—Charge controllers off at night. We turn the generator on before the SOC reaches 70%.

Solar— VDC Last daylight hours: (information obtained from magnum data)

Average 52.31

Minimum 48.40

Maximum 59.20

Notes on solar: Nothing to report

Propane Reading, station tank – 70 %

Propane Reading, director tank— 65 %

Propane Reading, intern tank— 55 %

Propane Reading, generator— 55 %

Ethanol Free Gasoline – 2 gallons

Water (Outpost tank) – 200 gallons

Sojourner rover used: Yes

Hours: 161.9

Beginning charge: 100

Ending charge: 100

Currently charging: Yes

Notes on rovers: Nothing to report

ATV’s Used: (Honda, 350.1, 350.2, 300): No

Reason for use: n/a

Oil Added? No.

ATV Fuel Used: 0 gallons

# Hours the ATVs were Used: 0

Notes on ATVs: Nothing to report

HabCar used and why, where? Yes, local

CrewCar used and why, where? Yes. Atila took it to drive to SLC to pick me up but lost control and went off the road close to I-70 on Highway 24. He was shaken and sore afterwards but the car was checked and appears to have no damage.

Luna used and why, where? No.

Campus wide inspection, if action taken, what and why? I have not completed my inspection fully but noted no damage/changes to the Robotic Observatory exterior, instrumentation nearby or the solar panels. Signs on the road and at the entrance have no damage/changes.

General notes and comments: I arrived back at MDRS on Wednesday night and Atila will be leaving MDRS on Saturday morning.

Summary of general operations:

Summary of internet: The original Hab internet service will be deactivated on March 1, 2022. Today the new system was to be installed, but the installer cancelled. I will be installing a Starlink system that hopefully can be used while we are waiting for HughesNet to get its act together.

Summary of suits and radios: Nothing to report

Summary of Hab operations: The toilet holding tank was cleaned for the built up waste that was making the tank fill twice as fast as when nominal. We were unable to see how successful we were, but will track the crew’s use to see if we cleared it completely. We will repeat this process in 4 weeks.

Summary of Outpost operations: Nothing to report

Summary of GreenHab operations: I inspected the GreenHab the morning after I arrived. From what few details I was getting from reports and photos, it was clear something was wrong. Nothing was thriving and plants were dying and underproducting. A more detailed inspection showed an unchecked infestation of white fly and contamination of soil by eggs and larva. I determined a collapse was near, salvage wasn’t possible, and we shut the building down. The crew removed all crops and consolidated the soil, and I will be finishing the removal of plant debris over the next few days. Heat is off and supplemental lighting turned off. This failure was the result of not one, but many small things. I was offsite and hadn’t been paying enough attention to the reports to realize that the crew was not reporting vital information about the condition of the plants. Our supplemental operations report had become truncated in a way that didn’t allow clear reporting on GreenHab issues by Atila. The GreenHab Officers and Atila didn’t have the knowledge and experience to recognize what was happening and I was offsite. Although we are requiring a general skill set for GreenHab Officer’s beginning next season, we do not have any requirements at the moment and it is one of two positions that are generally not filled by qualified people. Finally, but maybe not of much importance, I requested that the crew install a shade cloth weeks ago, and that was never gone. Dying plants were reported as drying plants. This is, however, why we do what we do—to learn from what mistakes we make. We probably won’t be able to produce anymore crops this season, because we need to sterilize the soil before we can replant and that requires heat, sunlight and time.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Robotic observatory is down. A part to repair it is due here Saturday.

Summary of health and safety issues: Nothing to report

Scientific Report – February 25th

Crew 263 Scientific Report 25-02-2022

SOL: 4

Name of person filing report: Léa Rouverand

Ultrasound surveillance:

On SOL 2, image capture of 5 organs for each Crew Members were performed using an ultrasound device and an Augmented Reality software. No issues were found regarding the hardware. Some issues in the protocol were identified and sent to experts at CNES (French Space Agency).

HF wave propagation experiment:

This experiment aims to study the propagation characteristics of the ionosphere for HF waves to achieve contact between MDRS and Toulouse (France). The Delta Loop Antennas were successfully set up on SOL 3.


Rotating platforms to study the effect of gravity on plant growth which were initially installed in the GreenHab were moved to the Science Dome on SOL 3.

Drone Astronaut Rescue:

The experiment’s aim is to assess the use of an ANAFI thermal drone for Astronaut Rescue. First trial was performed on SOL 4, the drone allowed to spot the Astronaut pair in 2 minutes and join them in 26 minutes. The use of the drone thermal feature is still to be investigated as it made the image more confusing. We also need to compare the rescue protocol to one without a drone; this is planned on SOL 16.

Atmospheric Experiment:

There is a total of 5 atmospheric experiments; LOAC (Light Optical Aerosol Counter), an Electrical Field Mill, Purple Air (Aerosol Counter), Mega-Ares (Electrical Field Counter) and a weather station. On SOL 4, the installation of 4 of them was finished. Mega-Ares still needs to be installed.

Augmented Reality Testing:

First trials of the Augmented Reality Application testing were performed between SOL 1 and SOL 4. Software issues with menu placement and hardware issues were identified.

On-going experiments: Sleep study using Dreem Headbands, TELEOP (analysis of spatial awareness and isolation on teleoperation tasks), Psychosocial and Cognitive questionnaires, Psychometric tests, Water quality monitoring. Nothing to declare on SOL 4.

Questions, concerns and requests to Mission Support: NaN

Crew 240 EVA Report February 8th

category eva-report]

Crew 240 EVA Report 08-02-2022

EVA # 9

Author: Julie Levita

Purpose of EVA: Exploration and looking out for interesting features for 3D mapping. Changing batteries on atmospheric experiment

Start time: 9:31

End time: 12:19

Narrative: Everything was fine, we did not have any issues changing batteries for the atmosphere experiment, and we saw beautiful landscapes at Moon Overlook. However, spacesuit batteries were abnormally low, so our engineer François has started trying to figure out the problem.

Destination: Moon Overlook, exploration zone D

Coordinates (use UTM NAD27 CONUS): 516 000, 425 4500

Participants: Julie Levita, Clément Plagne, Francois Vinet, Maxime Jalabert

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road. By foot : Brahe Hwy, Copernicus Hwy

Mode of travel: Rovers, foot

EVA Report February 1st

[category EVA-report]

Crew 240 EVA Report 02-01-2022

EVA #1

Author: Clément Plagne

Purpose of EVA: First in-sim EVA experience, acquaintance with procedures

Start time: 10:03

End time: 11:00

Narrative: Pre-EVA briefing went smoothly, and we drove to the spot safely and slowly. Finding my bearings with the map and getting my confidence back took a smidge more time than I expected, but we arrived at Marble Ritual without any delay. We walked on Marble Ritual, and a little bit around its rocky formation to get some bearings in the suits. Pierre felt confident enough in the suit to try and take out his drone, which I approved for a quick flight and a picture. A crew member mentioned that they were beginning to feel some back pain, and we therefore decided to call it for the walk and drive back to MDRS, which went well too. We were pleased to see that the batteries for both rovers stayed very high, as they had depleted fairly quickly on training day.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 518800 4250700

Participants: Clément Plagne, François Vinet, Pierre Fabre, Julie Levita

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road

Mode of travel: Rovers Spirit and Opportunity, foot.

Copyright © The Mars Society. All rights reserved. | Main Site